Algebraic Geometry and Integrable Systems related to String Theory

RIMS (Research Institute for Mathematical Sciences) Kyoto University

June, 12 – June, 16, 2000

Mon  Tues  Wed  Thurs  Fri
9:30–10:30 no talk  Ohta Eliashberg Morrison Morrison
10:50–11:50 no talk  Faber  Faber  Behrend Getzler
13:30–14:30 Fukaya  Ito  Morava  Bryan  no talk
14:50–15:50 Nakajima Sugiyama  TBA  Leung  no talk
16:10–17:10  Saito  Yoshioka  TBA  Takahashi no talk

June 12, (Monday)

• 13:30–14:30, Kenji Fukaya (Kyoto)

Floer homology for family

• 14:50–15:50, Hiraku Nakajima (Kyoto): Moduli of sheaves on blown up surfaces

• 16:10–17:10, Masa-Hiko Saito (Kobe)

Deformations of Okamoto–Painlevé pairs and Painlevé equations. ( Joint works with Takebe and Terajima )

June 13, (Tuesday)

• 9:30–10:30, Hiroshi Ohta (Nagoya):

Obstruction to and deformation of Lagrangian intersection Floer homology. (Joint works with Fukaya, Kontsevich, Oh, Ohta, Ono)

• 10:50–11:50 Carel Faber (Oklahoma State): Hodge integrals and Gromov-Witten theory I

• 13:30–14:30, Yukari Ito (Tokyo Metropolitan): McKay correspondence and T-duality

• 14:50–15:50, Ken-ichi Sugiyama (Chiba): Zeta functions of smooth manifolds and elliptic cohomology • 16:10–17:10, Kota Yoshioka (Kobe):

Brill-Noether problem for sheaves on K3 surfaces June 14, (Wednesday)

• 9:30–10:30, Y. Eliashberg, (RIMS/Stanford):

Few Examples from Symplectic Field Thoery

• 10:50–11:50, Carel Faber (Oklahoma State)

Hodge integrals and Gromov-Witten theory II

• 13:30–14:30, Jack Morava (The Johns Hopkins)

Topological gravity in dimensions two and four June 15, (Thursday)

• 9:30–10:30, David Morrison (Duke):

Theta-functions and six-manifolds ( after Hopkins and Singer)

• 10:50–11:50, Kai Behrend (RIMS/British Columbia)

Differential Graded Schemes

• 13:30–14:30, Jim Bryan (Tulane):

Multiple coverings, BPS states, and integrality in Gromov-Witten theory

• 14:50–15:50, Naichung Conan Leung (Minnesota):

Moduli of G-bundles over T^{4} and hyperkahler manifolds

• 16:10–17:10, Atushi Takahashi (RIMS)

Gromov–Witten Invariants and moduli of sheaves June 16, (Friday)

• 9:30–10:30, David Morrison (Duke):

Mirror symmetry and topology

• 10:50–11:50 Ezra Getzler (RIMS/Northwestern):

Higher genus Frobenius manifolds

All questions and comments on this conference should be sent to the organizer, Masa-Hiko Saito. Limited amount of support for participants inside Japan will be available. Please write to Saito before May 29.

Department of Mathematics, Faculty of Science Kobe University, Kobe, 657-8501, Japan mhsaito@math.kobe-u.ac.jp fax:+81-(0)78-803-5610 tel:+81-(0)78-803-5614